# Content-based image retrieval for multi-class volumetric radiology images: a benchmark study

### v1

@@ 2.4 Search and Retrieval
After creating the vector database, the search is performed using the embeddings extracted from slices of query volumes. The simplest way of retrieval is to match each 2D slice with the most similar 2D slice in the database. Here, we used cosine similarity. In Khun Jush et al. (2023) we introduced this method as the lower bound baseline for evaluating our proposed aggregation and sampling schemes. Similarly, in this work, we keep the slice-wise evaluation as the lower bound for the retrieval rate of our methods. We performed and evaluated image retrieval at the level of volumes and sub-volumes or regions. The difference between volume and region-based retrieval is as follows:
2.4.1 Volume-based
@@
For every slice within the query volume, the system retrieves the most similar slice from the database. Subsequently, the corresponding volume-id and its similarity score for each retrieved slice are stored in a hit-table similar to the hit-table shown in Figure 1. The aggregation method in Khun Jush et al. (2023) is a count-based method that retrieves per query volume the volume that has the most number of similar slices from the database Figure 1. Abacha et al. (2023) refers to this method as the evaluation based on frequency. For every retrieved slice, its corresponding volume-id is saved in the hit-table. The occurrence of volume-ids is then counted per each query volume. The volume with the highest count is selected as the most similar retrieved volume. Evaluation is then carried out based on the aggregated labels of the query volume and the most similar retrieved volume. This method retrieves the most similar volume per query volume. An overview is shown in Figure 2(a).
2.4.2 Region-based
The hit-table is generated following the process outlined in Section 2.4.1 with the same aggregation method. The key distinction lies in the fact that the subsequent retrieved volumes and final evaluation are predicated upon sub-volumes or anatomical regions. Meaning, that the occurrence of volume-ids is counted per each query region. The volumes with the highest count are selected as the most similar retrieved volumes. This means that in contrast to volume-based retrieval when all organs are under examination, multiple volumes are retrieved for each query volume. Naturally, only one organ can be searched (by querying only the slices that contain the limited organ view). The overview of this method is depicted in Figure 2(b).
Figure 2: Comparison of volume-based (a) and region-based (b) retrieval, in volume-based retrieval per each query volume one volume is retrieved while in region-based retrieval per each anatomical structure, one volume is retrieved.
2.5 Re-ranking
@@
Re-ranking in information retrieval involves the process of re-ordering the initially retrieved results to better align with the user’s information needs. This can be achieved through different methods such as relevance feedback, learning to rank algorithms, or incorporating contextual information (Ai et al., 2018; Guo et al., 2020; MacAvaney et al., 2019). Relevance feedback allows users to provide input on the initial results, which is then used to adjust the ranking (Ai et al., 2018). Learning to rank algorithms utilize machine learning techniques to re-rank results based on relevant features (Guo et al., 2020). Additionally, re-ranking methods may also consider contextual information such as user behavior, temporal relevance, or other relevant factors to better reflect the user’s current information needs, ultimately enhancing the overall quality of retrieved results (MacAvaney et al., 2019). A method based on contextualized information proposed in Khattab and Zaharia (2020) called ColBERT (Contextualized Late Interaction over BERT). ColBERT operates by generating contextualized representations of the query and the documents using BERT (Devlin et al., 2018). In this method, queries and documents are encoded into more detailed multi-vector representations, and relevance is gauged through comprehensive yet scalable interactions between these sets of vectors. ColBERT creates an embedding for each token in the query and document, and it measures relevance as the total of maximum similarities between each query vector and all vectors within the document (Santhanam et al., 2021). This late interaction approach allows for a more refined and contextually aware retrieval process, thereby enhancing the quality of information retrieval.
@@
Inspired by ColBERT we introduce a method in which filtering of the search space is performed and the total similarity of the entire target volume is considered to re-rank and score the retrieved volumes. To create an analogy to the ColBERT method each word can be considered as one slice and each passage of the database or each question of the query can be considered as one volume. Instead of the BERT encoder for the image retrieval task, the pre-trained vision models can be used to create the embeddings as discussed in Section 2.2.
@@
Figure 3: Overview of re-ranking: Step 1: Filtering based on at least one similar slice leads to the selection of candidate volumes Step 2: followed by similarity score computation using dot product on the normalized embedding matrices. Step 3: The final step involves max-pooling and summation to determine the top-scoring volumes for retrieval.
@@ 2.5.1 Step 1: Filtering
In the first step per each slice query, top k’ retrieval in the indexed space is performed based on cosine, L2 similarity, etc. In this step all the embedding are saved as vectors, one per each slice both for the query and the database. For each volume, that contained $N_{q}$ slices, maximum $N_{q}\times k^{\prime}$ volumes are retrieved. Thus, for $V$ query volumes or regions, in total maximum $N_{q}\times k^{\prime}\times V$ volumes are filtered. In the filtered volumes there is at least one slice that has a high similarity with its corresponding query volume or region. all the following operations are performed on the subsection of the filtered volumes and the re-ranking will be performed only on the filtered volumes.
@@ 2.5.2 Step 2: Similarity Score on Embedding Matrix
In this step, volumes are treated as matrices, where for each query $V_{q}$ there is a matrix of multiple embeddings. If the embeddings per slice have size $L$ and $N_{q}$ is the number of slices the loaded query matrix $M_{V_{q}}$ has size $Nq\times L$. Similarly, there is a matrix of multiple embeddings for each volume in the database. In this step, all the embeddings should be normalized such that the L2 norm is equal to 1. The result is that the dot-product of any two embeddings will be equivalent to their cosine similarity. The dot product of each query embedding matrix with size $Nq\times L$ to an embedding matrix of a volume in the database of size $M_{v_{x}}\times L$ results in a similarity score matrix of size $N_{q}\times M_{v_{x}}$ where $x=1,2,...$ denoting volume-id number.
@@ 2.5.3 Step 3: Final Score and Re-ranking
To compute the score of each volume, its dimension across the volume is reduced via max-pooling (i.e. representing the most similar slice in the target volume for each query slice). Across the query dimension, a summation is performed (i.e. representing the total score of the similarity of all the slices of the query to the whole volume in the database). Finally, the k’ documents are sorted by their total scores, and the volume/volumes with the maximum overall score are retrieved.
@@ 3 Evaluation
In this section, we evaluate the retrieval recall of the methods explained in Section 2.4 and Section 2.5. Since finding extra anatomical regions is not critical for this study (that would be the anatomical regions that are present in the retrieved volume/slice and are missing from the query volume/slice, i.e. false positives (FP)) we do not discuss the precision metric. The results are presented for 29 anatomical structures presented in LABEL:tab:_mapping_table and 104 anatomical structures that were originally presented in Wasserthal et al. (2023). In the tables presented in this section, the average and standard deviation (STD) columns are aimed to highlight difficult classes across models (low average) and the ones that have higher variations among models (higher STD). The average and STD rows show the average and STD over all the classes for each model.
@@ 3.1.1 Slice-wise
In the computation of slice-wise recall, per each slice, if the retrieved slice contains the same anatomical region/regions the corresponding class/classes are considered as the true positive class (TP). If the query slice contains anatomical regions that are not present in the retrieved slice that class is considered a false negative (FN).
LABEL:tab:slicewise-recall-28-regions and LABEL:tab:slicewise-recall-104-regions show the retrieval recall of 29 coarse anatomical regions and 104 original TS anatomical regions, respectively, using the slice-wise method (lower bound).
In slice-wise retrieval, DreamSim is the best-performing model with retrieval recall of .$849\pm.148$ and .$9.7$129 for coarse and original TS classes, respectively. ResNet50 pre-trained on fractal images has the lowest retrieval recall almost on every anatomical region for 29 and 104 classes. This is however expected due to the nature of synthetic generated images.
@@ 3.1.2 Volume-based
This section presents the recall of volume-based retrieval explained in Section 2.4.1 An overview of the evaluation is shown in Figure 2(a). In volume-based retrieval, per each query volume, one volume is retrieved. In the recall computation, the classes present in both the query and the retrieved volume are considered TP classes. The classes that are present in the query volume and are missing from the retrieved volume are considered FN.
LABEL:tab:volumewise-recall-29-regions and LABEL:tab:volumewise-recall-104-regions present the retrieval recall of the volume-based method on 29 and 104 classes, respectively. The overall recall rates are increased compared slice-wise which is expected due to the aggregation and effects of neighboring slices.
@@ 3.1.3 Region-based
This section presents the recall of region-based retrieval. An overview of the evaluation is shown in Figure 2(b). In region-based retrieval, per each anatomical region in the query volume, one volume is retrieved. In the recall computation, the classes present in both the sub-volume of the query and the corresponding retrieved volume are considered TP classes. The classes that are present in the query sub-volume and are missing from the retrieved volume are considered FN.
LABEL:tab:regionbased-recall-29-regions and LABEL:tab:regionbased-recall-104-regions present the retrieval recalls. Compared to volume-based retrieval the average retrieval for the regions is higher. The performance of the models is very close. DreamSim performs slightly better with an average recall of .$7.9$037 for coarse anatomical regions and .$983\pm.032$ for 104 anatomical regions. The retrieval recall for many classes is $1.0$. The standard deviation among classes and the models is low, with the highest standard deviation of .05.05.05.05 and .092.092.092.092, respectively.

### v2

@@ 2.4 Search and Retrieval
After creating the vector database, the search is performed using the embeddings extracted from slices of query volumes. The simplest way of retrieval is to match a 2D query slice $q$ with the most similar 2D slice in the database $s^{*}$ by finding the slice-embedding that maximizes the cosine similarity with respect to the embedding associated with $q$, i.e.
$s^{*}=\underset{s\in\text{Data}}{\mathrm{argmax}}\frac{\langle\phi(s),\phi(q)\rangle}{\|\phi(s)\|_{2}\|\phi(q)\|_{2}}=\underset{s\in\text{Data}}{\mathrm{argmax}}\left\langle v_{s},\frac{\phi(q)}{\|\phi(q)\|_{2}}\right\rangle$ (1)
[…]
2.4.1 Volume-based retrieval and evaluation
Figure 2: Volume-based retrieval: For a query volume $V_{q}$ covering a range of anatomical regions, a volume is retrieved that should cover the same anatomical regions. The similarity search is based on all slices from the query volume.
@@ 2.4.1 Volume-based retrieval and evaluation
In this setting, we want to assess the capability of the search system to retrieve an image volume that covers the same anatomical regions overall as a given query volume $V_{Q}=[q_{1},...,q_{n}]$. For every slice $q_{i}$ from the query volume $V_{Q}$, the system retrieves the most similar slice $s_{i}^{*}$ from the database using (1). Subsequently, the corresponding volume-id and its similarity score are stored in a hit-table similar to the hit-table shown in Figure 1. We then apply the count-based aggregation scheme from Khun Jush et al. (2023) that uses the hit-table to identify the volume $V_{R}$ with the most hits given the query volume at hand, see Figure 1. The system then returns $V_{R}$ as the final result of the similarity search given the query volume $V_{Q}$. Abacha et al. (2023) refers to this method as the retrieval based on frequency. The evaluation of the search system on the proposed benchmark dataset is done using the measure of Recall (True Positive Rate). To this end, the aggregated labels of the query volumes are compared against the aggregated labels of the associated retrieved volumes (see 3.1.2). The recall in this setting is high if the aggregated query and retrieved volume labels coincide. An overview is shown in Figure 2.
2.4.2 Region-based retrieval and evaluation
In this setting the search system is queried with an image (sub-)volume which is constrained to a specific anatomical sub-region (e.g. liver, pancreas, heart,…). For each anatomical region, we want to individually assess the capability of the system to retrieve an image volume containing the anatomical region.
The query (sub-)volumes for different anatomical regions are generated as follows. Given a selected anatomical region $r$ and a query image volume $V_{Q}=[q_{1},...,q_{n}]$, the smallest subset slices $V_{Q,r}=[q_{m},...,q_{k}]\subset V_{Q}$ is chosen that entirely contains the anatomical region $r$ visible in $V_{Q}$. Based on the sub-volume $V_{Q,r}$ a similarity search is conducted to build up a hit-table, and the count-based aggregation is conducted to finally retrieve for this query the volume with most hits, as described in Section 2.4.1.
Figure 3: Region-based retrieval. Anatomical regions are considered individually. A sub-volume constrained to an anatomical region of interest $r$ is generated and fed to the search system to retrieve a volume containing the anatomical region. A case is considered a True Positive (TP) if the retrieved case contains the region $r$ at some location.
[…]
2.4.3 Localized retrieval and evaluation
Figure 4: Localized retrieval. Anatomical regions are considered individually. A sub-volume constrained to the anatomical region of interest $r$ is generated and fed to the search system to retrieve a volume containing the same anatomical region. A case is only considered as True Positive (TP) if at least one of the slices in the retrieved volume contains the region $r$.
In this setting, the system is queried with an image sub-volume which is constrained to a specific anatomical sub-region (e.g. liver, pancreas, heart,…). For each anatomical region, we want to individually assess the capability of the system to retrieve an image volume containing the anatomical region and to localize the region of interest within the retrieved volume.
[…]
Another measure to assess the capability of the system to localize a region can be defined as the ratio of the slices that actually contain the subregion $r$ in the retrieved volume to the total number of slices hit in the retrieved volume. In detail, the localization-ratio (LR) is defined as:
LR=|slices hit in retrieved vol. that contain r||slices hit in retrieved $\mathrm{LR}=\frac{|\text{slices hit in retrieved vol. that contain r}|}{|\text{slices hit in retrieved vol.}|}$ (2)
For example, the query consists of $60$ slices related to region $r$. The table representing the top 3 volumes hit count is $[48,21,4]$. In the volume with the top hit count, 12 out of the 48 hit slices actually contain region $r$, indicating successful localization. The localization-ratio is then given by $12/48=.25$.
2.5 Re-ranking retrieval and evaluation
@@
Re-ranking in information retrieval involves the process of re-ordering the initially retrieved results to better align with the user’s information needs. This can be achieved through different methods such as relevance feedback, learning to rank algorithms, or incorporating contextual information (Ai et al., 2018; Guo et al., 2020; MacAvaney et al., 2019). Relevance feedback allows users to provide input on the initial results, which is then used to adjust the ranking (Ai et al., 2018). Learning to rank algorithms utilizes machine learning techniques to re-rank results based on relevant features (Guo et al., 2020). Additionally, re-ranking methods may also consider contextual information such as user behavior, temporal relevance, or other relevant factors to better reflect the user’s current information needs, ultimately enhancing the overall quality of retrieved results (MacAvaney et al., 2019). A method based on contextualized information proposed in Khattab and Zaharia (2020) called ColBERT (Contextualized Late Interaction over BERT). ColBERT operates by generating contextualized representations of the query and the documents using BERT (Devlin et al., 2018). In this method, queries and documents are encoded into more detailed multi-vector representations, and relevance is gauged through comprehensive yet scalable interactions between these sets of vectors. ColBERT creates an embedding for each token in the query and document, and it measures relevance as the total of maximum similarities between each query vector and all vectors within the document (Santhanam et al., 2021). This late interaction approach allows for a more refined and contextually aware retrieval process, thereby enhancing the quality of information retrieval.
@@
Inspired by ColBERT we introduce a two-stage method in which filtering of the search space is performed and the total similarity of the entire target volume is considered to re-rank and score the retrieved volumes. To create an analogy to the ColBERT method each word can be considered as one slice and each passage of the database or each question of the query can be considered as one volume. Instead of the BERT encoder for the image retrieval task, the pre-trained vision models can be used to create the embeddings as discussed in Section 2.2.
@@
Figure 5: Overview of re-ranking. Step 1: Filtering based on at least one similar slice leads to the selection of candidate volumes Step 2: followed by similarity score computation using dot product on the normalized embedding matrices. Step 3: The final step involves max-pooling and summation to determine the top-scoring volumes for retrieval.
@@ 2.5.1 Step 1: Filtering
We assume that the embeddings have been computed and stored as vectors for all image data. In the first step, for each slice $q_{i}$ of the query volume $V_{Q}=[q_{1},...,q_{N}]$, the most similar slice $s_{i}^{*}$ from the database is retrieved via similarity search (1). We denote volumes associated to $s_{1}^{*},...,s_{N}^{*}$ by $V_{1},...,V_{N}$. Clearly, for some indices $i,j$ the retrieved sliced can belong to the same volume, i.e. $V_{i}=V_{j}$. We filter out duplicates for our list and let $V_{1},..,V_{M}$ with $M\leq N$ denote the unique volumes associated with the retrieved slices. These unique volumes resulting from this initial search stage are considered as candidates for a more compute-intensive re-ranking procedure which takes all similarity interactions between all the volume slices into account.
@@ 2.5.2 Step 2: Similarity Score on Embedding Matrix
We assume that our slice embedding $\phi$ maps each slice to a vector of fixed length $L$. Hence, for any image volume $V=[v_{1},...,v_{n}]$ with $n$ slices $v_{1},...,v_{n}$ we can compute the $L_{2}$-normalized embedding matrix
$M_{V}=\left[\frac{\phi(v_{1})}{\|\phi(v_{1})\|_{2}},...,\frac{\phi(v_{n})}{\|\phi(v_{n})\|_{2}}\right]$ (3)
of dimension $n\times L$. Given another image volume $W=[w_{1},...,w_{m}]$ of size $m\times L$ we can compute the similarity matrix of size $n\times m$ via the matrix dot product:
$\mathrm{SIM}(M_{V},M_{W})=\left[\frac{\langle\phi(v_{i}),\phi(w_{j})\rangle}{\|\phi(v_{i})\|_{2}~{}\|\phi(w_{j})\|_{2}}\right]_{\begin{subarray}{c}i=1,...,n\\ j=1,...,m\end{subarray}}$ (4)
The entry $(i,j)$ of this matrix contains the cosine-similarity score of the embeddings related to slice $i$ of volume $V$ and slice $j$ of volume $W$. For our re-ranking, we first compute in this step, the embedding matrix $M_{V_{Q}}$ for the query volume $V_{Q}$ and the embedding matrices $M_{V_{1}},...,M_{V_{M}}$ for all unique retrieved volumes $V_{1},...,V_{M}$ from Step 1. Then we compute the similarity matrices $\mathrm{SIM}(M_{V_{Q}},M_{V_{k}})$ for $k=1,...,M$.
@@ 2.5.3 Step 3: Final Score and Re-ranking
To compute the rank score of each volume $V_{i}$, the associated similarity matrix $\mathrm{SIM}(M_{V_{Q}},M_{V_{i}})$ is max-pooled row-wise which corresponds to determine the slice in $V_{i}$ with highest cosine similarity to a given slice in $V_{Q}$. The resulting vector of length $n$ is then summed to obtain the total maximum slice similarity as final rank score (RS), i.e. for $k=1,...,M$ we compute:
$\mathrm{RS}(V_{k})=\sum_{i=1}^{n}\max_{j=1,...,m_{k}}\mathrm{SIM}(M_{V_{Q}},M_{V_{k}})_{i,j}$ (5)
where ${V_{k}}_{j}$ denotes the $j$-th slice of volume $V_{k}$ and $m_{k}$ is the total number of slices of ${V_{k}}$. The retrieved volumes are re-ranked according to their RS, i.e. the volume with the highest score is considered as the most relevant search result.
2.5.4 Localization with Re-ranking
In section 2.4.3 we proposed a measure for assessing the capabilities of the search system to localize anatomical regions of interest. This localization measure relied on the hit slices that were finally retrieved by count-based aggregation. As described above, re-ranking may select a volume with a lower total hit count as the final search result based on a more fine-grained similarity measure. We propose to utilize this similarity information to identify slices of maximal similarity with the query data in order to localize the anatomical region of interest.
In detail, after having obtained the most relevant search result via re-ranking, we need to identify the most relevant slices that could contain the anatomical region of interest $r$. To this end, we consider the slices that highly contribute to the final RS score (5). Since the full similarity matrices (4) for all candidates from the first search stage are available, we can utilize these to identify the locations of maximum similarity between the query and candidate slices.
Assume that the volume $V_{k^{\star}}$ obtained the highest RS score in the re-ranking process and therefore was returned as the final result after re-ranking, i.e. $k^{*}=argmax_{k}(RS(V_{k}))$. Further assume that $V_{k^{\star}}$ consists of $m_{k^{\star}}$ slices. First, we compute the vector $m_{\mathrm{SIM}}$ of size $m_{k^{\star}}$ containing for each slice of $V_{k^{\star}}$ the maximum similarity with the query slices:
$m_{\mathrm{SIM}}(V_{k^{\star}})=\left[\max_{i=1,...,n}\mathrm{SIM}(M_{V_{Q}},M_{V_{k^{\star}}})_{i,j}\right]_{j=1,...,m_{k^{\star}}}$ (6)
Hence, to determine the $L\leq m_{k^{\star}}$ slices of $V_{k^{\star}}$ with the maximum similarity, we determine the indices of the $L$ largest values of the vector $m_{\mathrm{SIM}}$. These slices are considered candidate locations for the anatomical region $r$. In our experiments, we chose $L=15$. The localization-ratio (2) can be adopted in the case of re-ranking by determining how many of the $L$ most similar slices actually contained the anatomical region of interest.
LR=|slices among the L most similar slices of Vk that contain $\mathrm{LR}=\frac{|\text{slices among the $L$ most similar slices of $V_{k^{\star}}$ that contain r}|}{L}$ (7)
@@ 3 Evaluation
In this section, we evaluate the retrieval recall of the methods explained in Section 2.4 and Section 2.5. The results related to the 29 coarse anatomical structures from LABEL:tab:_mapping_table and the results related to the original 104 fine-grained anatomical structures fromWasserthal et al. (2023) are presented separately in the following. In the tables presented in this section, the average and standard deviation (STD) columns allow identifying difficult classes across models (low average) and the ones that have higher variations among models (higher STD). The average and STD rows show the average and STD over all the classes for each model.
@@ 3.1.1 Slice-wise
Detailed computation of the recall measure for different retrieval methods is explained in Section 2.4. LABEL:tab:slicewise-recall-28-regions and LABEL:tab:slicewise-recall-104-regions show the retrieval recall of 29 coarse anatomical regions and 104 original TS anatomical regions, respectively, using the slice-wise method. The slice-wise recall is considered the lower bound recall because for a perfect recall all the anatomical regions present in the query slice should appear in the retrieved slice.
In slice-wise retrieval, DreamSim is the best-performing model with retrieval recall of .$863\pm.107$ and .$9.7$129 for coarse and original TS classes, respectively. ResNet50 pre-trained on fractal images has the lowest retrieval recall almost on every anatomical region for 29 and 104 classes. This is however expected due to the nature of synthetic generated images.
@@ 3.1.2 Volume-based
This section presents the recall of volume-based retrieval explained in Section 2.4.1 An overview of the evaluation is shown in Figure 2. In volume-based retrieval, per each query volume, one volume is retrieved. In the recall computation, the classes present in both the query and the retrieved volume are considered TP classes. The classes that are present in the query volume and are missing from the retrieved volume are considered FN.
LABEL:tab:volumewise-recall-29-regions and LABEL:tab:volumewise-recall-104-regions present the retrieval recall of the volume-based method on 29 and 104 classes, respectively. The overall recall rates are increased compared to slice-wise retrieval which is expected due to the aggregation and contextual effects of neighboring slices.
@@ 3.1.3 Region-based
This section presents the recall of region-based retrieval. An overview of the evaluation is shown in Figure 3. In region-based retrieval, per each anatomical region in the query volume, one volume is retrieved. In the recall computation, the classes present in both the sub-volume of the query and the corresponding retrieved volume are considered TP classes. The classes that are present in the query sub-volume and are missing from the retrieved volume are considered FN.
LABEL:tab:regionbased-recall-29-regions and LABEL:tab:regionbased-recall-104-regions present the retrieval recalls. Compared to volume-based retrieval the average retrieval for the regions is higher. The performance of the models is very close. DreamSim performs slightly better with an average recall of .$7.9$037 for coarse anatomical regions and .$983\pm.032$ for 104 anatomical regions. The retrieval recall for many classes is $1.0$. The standard deviation among classes and the models is low, with the highest standard deviation of .076.076.076.076 and .092.092.092.092, for coarse and fine classes respectively.